Graph f(x) = 3x + 2 explain in wordthe points to graph.
Answer by fractalier(6550)

To graph f(x) = 3x + 2 on a coordinate plane, merely choose any arbitrary value for x (usually 2 or 1 or 0 or -1 or -2 work), and see what value the function has for that value of x.
Thus choose x = 1.
f(1) = 3(1) + 2 = 5.
Therefore, our first point to graph is (1, 5)
Then choose x = 2.
here f(2) = 3(2) + 2 = 8.
Our second point is (2, 8) and so on...